Jesse and her brothers Nick and Owen are saving money over the summer. Each week, Jesse saves twice as much as Owen. Owen saves $5 more than nick. At the end of four weeks, the three of them have saved a total of $220. How much money does each person save per week?
Found 2 solutions by garettbuff, MathTherapy:
Answer by garettbuff(18)

Jesse=8(x+5)
Owen=4(x+5)
Nick=4x
So 16x+60=220
16x=120
4x=30
X=30/4
Plug in x to the names

Answer by MathTherapy(10552)


Jesse and her brothers Nick and Owen are saving money over the summer. Each week, Jesse saves twice as much as Owen. Owen saves $5 more than nick. At the end of four weeks, the three of them have saved a total of $220. How much money does each person save per week?
Amount saved each week by: